a ball is thrown into the air in such a way that its distance(d), in feet, above the ground at any time(t), in seconds, is given by the equation d=40t-5t^2.
how many seconds will it take for the ball to reach a height of 80ft above the ground?

Answer by checkley77(12844)

d=40t-5t^2.
80=40T-5T^2
5T^2-40T+80=0
5(T^2-8T+16)=0
5(T-4)(T-4)=0
T-4=0
T=4 SEC. ANS.
PROOF:
80=40*4-5*4^2
80=160-5*8
80=160-80
80=80